OpenStudy (anonymous):

Find f(x) and g(x) so that the function can be described as y = f(g(x)). y =8/x^2 + 2

OpenStudy (anonymous):

well one way to do this is to say that \[f(x)= \frac{8}{x} + 2\] and \[g(x)=x^{2}\]

OpenStudy (zzr0ck3r):

f(x) = 8/(x^2+2) , g(x) = x

OpenStudy (zzr0ck3r):

:P

OpenStudy (zzr0ck3r):

or f(x) = 8/x^2 + 2 g(x) = x

OpenStudy (anonymous):

thanks! :)

OpenStudy (zzr0ck3r):

your teacher will love that:)

OpenStudy (anonymous):

Isn't there a million ways you could do this? It's just most fun to do the simpliest?

OpenStudy (zzr0ck3r):

I think mine is the simplest

OpenStudy (zzr0ck3r):

its sort of a joke but it works

OpenStudy (anonymous):

Nothing like being a smartass mathematically!